OpenStudy (anonymous):

Explain the steps that are necessary to solve 4x - 2(x + 3) = x + 1

OpenStudy (anonymous):

distribute the 2 , combine like terms solve for x

OpenStudy (anonymous):

4x-2(x+3) = x+1 4x-2x-6=x+1 2x-x=1+6 x=7